# Degradation-Aware All-in-One Image Restoration via Latent Prior Encoding

### v2

@@ Abstract
Real-world images often suffer from spatially diverse degradations such as haze, rain, snow, and low-light, significantly impacting visual quality and downstream vision tasks. Existing all-in-one restoration (AIR) approaches either depend on external text prompts or embed hand-crafted architectural priors (e.g., frequency heuristics); both impose discrete, brittle assumptions that weaken generalization to unseen or mixed degradations. To address this limitation, we propose to reframe AIR as learned latent prior inference, where degradation-aware representations are automatically inferred from the input without explicit task cues. Based on latent priors, we formulate AIR as a structured reasoning paradigm: (1) which features to route (adaptive feature selection), (2) where to restore (spatial localization), and (3) what to restore (degradation semantics). We design a lightweight decoding module that efficiently leverages these latent encoded cues for spatially-adaptive restoration. Extensive experiments across six common degradation tasks, five compound settings, and previously unseen degradations demonstrate that our method outperforms state-of-the-art (SOTA) approaches, achieving an average PSNR improvement of 1.68 dB while being three times more efficient.
@@ 1 Introduction
Real-world images are frequently degraded by factors like haze, rain, snow, low-light, and motion blur Tian et al. (2025a); Jiang et al. (2025). These spatially varying degradations may occur alternately or even overlap, depending on environmental and capture conditions Guo et al. (2024). Such entangled degradation patterns significantly reduce visual quality, resulting in impaired downstream vision tasks Tian et al. (2025a); Jiang et al. (2025). Recent deep learning-based single-task methods Dong et al. (2020); Valanarasu et al. (2022); Chen et al. (2021); Cai et al. (2023); Wang et al. (2022); Zhang et al. (2017) have made significant progress in addressing individual restoration challenges. However, deploying separate task-specific (TS) networks for each degradation type is computationally expensive and impractical, driving interest in AIR frameworks Jiang et al. (2025).
@@
Current AIR approaches can be broadly categorized into two paradigms. (1) Adaptive feature learning methods, such as ADAIR Cui et al. (2024) and AirNet Li et al. (2022), utilize hand-crafted frequency priors or architectural inductive biases to automatically differentiate degradation types. Although these methods circumvent manual specification, they inherently depend on pre-defined assumptions about degradation characteristics, which often fail to generalize to novel or compound corruptions Gao et al. (2024). (2) Prompt-based restoration methods, such as PromptIR Vaishnav et al. (2023) and DFPIR Tian et al. (2025a), rely on explicit, manually provided prompts to guide the restoration network by specifying degradation types. These methods offer flexible control but face a fundamental ”chicken-and-egg” dilemma: in real-world scenarios, the degradation type and location are rarely known beforehand, yet the network requires this information to perform effective restoration Jiang et al. (2025). Moreover, TS instructions constrain generalization, particularly in cases of mixed degradations or varying homogeneous degradations (e.g., different noise levels). As illustrated in Fig. 1, both existing AIR paradigms have limited robustness and practical applicability in unconstrained environments.
@@
To overcome these limitations, we reframe AIR as a latent prior inference problem. Unlike existing methods Vaishnav et al. (2023); Cui et al. (2024); Tian et al. (2025a) that rely on explicit degradation prompting or predefined architectural inductive biases, we propose learning degradation-aware representations directly from the degraded image through a multi-level feature descriptor inspired by the variational autoencoder (VAE) Kingma & Welling (2013). Our learned prior eliminates the need for manual degradation hints and enables spatially adaptive restoration of diverse and unseen degradations. Guided by these learned priors, the proposed degradation-aware AIR (DAIR) framework incorporates a ”where, which, what” reasoning paradigm: it learns degradation-aware feature selection (which), localizes corrupted regions using spatial attention maps (where), and adaptively fuses multi-scale global representations (what). Our unified strategy substantially enhances the flexibility and generalizability of blind restoration methods.
@@
  • DAIR consistently outperforms SOTA on six common restoration tasks (e.g., snow, low-light), five compound degradations (e.g., haze + rain, low-light + haze + rain), and unseen degradations, achieving an average 1.68 dB PSNR gain. It also improves downstream tasks on images with unknown multi-type degradations, e.g., boosting YOLOv12-L Tian et al. (2025b) object detection by up to 7.40 mAP over SOTA AIR methods, highlighting strong generalizability (Fig. 1).
@@ 2.1 Image-to-image Translation
Early image restoration methods were designed to address individual TS settings, such as denoising Zhang et al. (2017); Pang et al. (2021); Zhang et al. (2018a), dehazing Dong et al. (2020); Qin et al. (2020), derainingJiang et al. (2020); Ren et al. (2019), low-light image enhancement (LLIE) Wei et al. (2018); Yi et al. (2023), or deblurring Cho et al. (2021); Nah et al. (2017). These task-dependent approaches achieved impressive results; their limited scope prevented generalization across diverse restoration challenges. Transformer-based models, such as Restormer Zamir et al. (2022) and Uformer Wang et al. (2022), have explored multiple degradation scenarios; however, they still require separate training for individual tasks. To address this, recent methods shift towards AIR, aiming to handle diverse degradations within a unified framework. AirNet Li et al. (2022) pioneered this field by introducing contrastive learning to extract degradation representations, guiding restoration for unknown corruptions. ADAIR Cui et al. (2024) recalibrate features using adaptive frequency statistics, enabling task-agnostic restoration. Recent methods, such as DiffUIR Zheng et al. (2024), have also explored generative techniques (i.e., latent diffusion Rombach et al. (2022)). These unified approaches mark a significant step forward, minimizing TS training and advancing toward generalizable AIR solutions.
@@ 2.2 Prompt-Guided All-in-one Restoration
Image-to-image translation-based method commonly fails in separating unique degradations Brooks et al. (2023); Jiang et al. (2025). To counter this, recent AIR Gao et al. (2024); Vaishnav et al. (2023); Tian et al. (2025a) methods have incorporated MP to provide prior knowledge of degradation cues. PromptIR Vaishnav et al. (2023) integrates text-based prompts into a transformer framework, leveraging degradation-specific semantic information to guide the restoration process. OneRestore Guo et al. (2024) introduced visual and text prompts within a transformer-based framework to encode TS information. Recently, DFPIR Tian et al. (2025a) introduced a degradation-aware feature perturbation method that utilizes CLIP-encoded Radford et al. (2021) text prompts to guide channel shuffling and attention masking, enabling unified restoration across diverse degradations. UHD-Processor Liu et al. (2025) introduced a VAE-based framework with progressive frequency learning and MP for ultra-high-definition image restoration. Notably, unlike UHD-Processor and other prompt-based methods, our method learns latent degradation representations directly from the input, enabling blind restoration without user guidance.
Figure 2: Overview of the proposed DAIR framework: (a) Overall architecture, (b) Latent prior encoding (which), (c) Degradation mapping (where), (d) Latent fusion (what), (e) 3WD restoration module. Best viewed with Zoom.
@@ 3 Method
Fig. 2 illustrates the details of our framework. Given a degraded input image $\mathbf{I}_{deg}\in\mathbb{R}^{3\times H\times W}$ with unknown degradation $\mathcal{D}_{R}$, our framework learn to infer stage-wise latent codes $x\ell$ ($\ell=0,\dots,3$) and a bottleneck global descriptor $\mu$, serving as degradation priors. The encoder employs a two-branch architecture that separately processes multi-scale structural features (e.g., luminance $L_{\ell}$) and color features (e.g., chrominance $C_{\ell}$), enabling more effective representation of both spatial and chromatic information Sharif et al. (2025); Yan et al. (2025). At each stage, $L_{\ell}$ and $C_{\ell}$ interact with $x_{\ell}$ via Degradation-Aware Encoder Blocks (DAEB), leveraging degradation information into the feature extraction process. The branches are fused as $F_{\ell}=L_{\ell}+C_{\ell}$ and used as residual refined feature propagation to the decoder. Concurrently, a mapping block transforms $(L_{\ell},C_{\ell},x_{\ell})$ into a compact DM $g_{\ell}$, which is also forwarded to the decoder for explicit degradation localization guidance. At the bottleneck, a $\mu$-guided fusion block merges the deepest LC features with the global degradation latent using adaptive scaling and shifting Perez et al. (2018), resulting in $F_{\mu}=\gamma\odot[L_{3}+C_{3}]+\beta$. The decoder reconstructs the clean image from $F_{\mu}$, leveraging the degradation focus map $g_{\ell}$ and the selected features $F_{\ell}$ derived from latent priors. This architecture enables the model to automatically identify and restore a wide range of degradations without explicit user input or task specification.
@@ 3.1 Learning Degradation Reasoning with Latent Prior
We propose a multi-scale latent descriptor Kingma & Welling (2013) that directly infers continuous degradation representations from corrupted images without user-guided MP. Given $\mathbf{I}_{deg}\in\mathbb{R}^{3\times H\times W}$, the encoder extracts hierarchical feature maps $\{\mathbf{x}_{\ell}\}_{\ell=0}^{3}$ using convolutional layers and residual channel-attention Hu et al. (2018) blocks that emphasize degradation-relevant cues. At the deepest stage, multi-head self-attention Vaswani et al. (2017) refines the features to obtain the bottleneck representation $\mathbf{z}$. The VAE parameterizes a distribution with $\bm{\mu}=f_{\mu}(\mathbf{z})$ and $\log\bm{\sigma}^{2}=f_{\sigma}(\mathbf{z})$, sampling $\mathbf{z}_{reparam}=\bm{\mu}+\bm{\sigma}\odot\bm{\epsilon}$ where $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. The decoder reconstructs the input image from the sampled latent. This multi-objective training ensures that hierarchical features $\{\mathbf{x}_{\ell}\}$ capture degradation-specific representations at multiple scales, while bottleneck statistics $\bm{\mu}$ encode global degradation semantics. Both serve as degradation-aware latent cues to condition the restoration model:
@@ 3.1.1 Latent Prior Encoding
Unlike existing methods (PromptIR Vaishnav et al. (2023), DFPIR Tian et al. (2025a), CAPTNet Gao et al. (2024)) that inject prompts during decoding, we inject VAE-derived degradation priors during encoding. This enables our encoders to generate both degradation-aware and degradation-agnostic skip connections, providing richer guidance for reconstruction. We separately encode the luminance and chrominance (LC) components Smith (1978)($\mathbf{L}_{\text{deg}}$, $\mathbf{C}_{\text{deg}}$) of the degraded input using dedicated LC encoders Yan et al. (2025); Sharif et al. (2025) for better structural and chromatic information. Fig. 2 (b) illustrates our latent prior encoding. Two parallel encoders, $\mathcal{E}_{L}(\cdot)$ and $\mathcal{E}_{C}(\cdot)$, process these components across multiple scales using DAEB. At stage $\ell$, we injects latent degradation priors $\mathbf{x}_{\ell}$ via value modulation:
@@ 3.1.2 Learnable Degradation Map
Traditional skip connections propagate noise from degraded inputs to the decoder, without spatial awareness of the degradation Mao et al. (2016). We counter these limitations with learnable DM that encode spatially-varying degradation patterns for effective reconstruction (Fig. 2 (c)). At encoder stage $\ell$, our Degradation Mapping Block generates scale-specific maps $g_{\ell}$ from LC features $L_{\ell},C_{\ell}$ and latent prior encoding $\mathbf{x}_{\ell}$. We first fuse complementary LC information: $\mathbf{f}^{LC}_{\ell}=\text{Conv}_{3\times 3}(L_{\ell}+C_{\ell})$. To capture frequency-domain degradation characteristics, we enhance features with spectral information via Fast Fourier Transform Cooley & Tukey (1965): $\mathcal{F}(L_{\ell})=\text{FFT}(L_{\ell}-\mu(L_{\ell}))$, $\mathcal{F}(C_{\ell})=\text{FFT}(C_{\ell}-\mu(C_{\ell}))$. The magnitude and phase components are concatenated and projected:
@@ 3.1.3 Latent Encoded Fusion
At the encoder bottleneck, we leverage the deepest LC features $\mathbf{L}_{3},\mathbf{C}_{3}\in\mathbb{R}^{B\times 320\times\frac{H}{8}\times\frac{W}{8}}$, along with the latent global descriptor $\mu$. As shown in Fig. 2 (d), our fusion mechanism integrates three components: (i) intra-branch self-attention Sharif et al. (2025); Cai et al. (2023) for independent modality refinement, where $\hat{\mathbf{L}}_{3}=\text{MHSA}(\mathbf{L}_{3})$ and $\hat{\mathbf{C}}_{3}=\text{MHSA}(\mathbf{C}_{3})$; (ii) cross-branch attention Sharif et al. (2025) for complementary information exchange, where chrominance attends to luminance: $\mathbf{F}=\text{LayerNorm}\left(\text{MHCA}(\hat{\mathbf{C}}_{3},\hat{\mathbf{L}}_{3})+\hat{\mathbf{C}}_{3}+\hat{\mathbf{L}}_{3}\right)$, enabling the model to leverage structural and color information while preserving individual branch characteristics; and (iii) adaptive feature modulation conditioned on the global degradation prior $\mu$, combining LC features through adaptive scaling and shifting for TS feature modulation while maintaining underlying image characteristics. Unlike standard FiLM Perez et al. (2018) that applies uniform modulation, our approach makes degradation-aware reconstruction:

### v3

@@ Abstract
Real-world images often suffer from spatially diverse degradations such as haze, rain, snow, and low-light, significantly impacting visual quality and downstream vision tasks. Existing all-in-one restoration (AIR) approaches either depend on external text prompts or embed hand-crafted architectural priors (e.g., frequency heuristics); both impose discrete, brittle assumptions that weaken generalization to unseen or mixed degradations. To address this limitation, we propose to reframe AIR as learned latent prior inference, where degradation-aware representations are automatically inferred from the input without explicit task cues. Based on latent priors, we formulate AIR as a structured reasoning paradigm: (1) which features to route (adaptive feature selection), (2) where to restore (spatial localization), and (3) what to restore (degradation semantics). We design a lightweight decoding module that efficiently leverages these latent encoded cues for spatially-adaptive restoration. Extensive experiments across six common degradation tasks, five compound settings, and previously unseen degradations demonstrate that our method outperforms state-of-the-art (SOTA) approaches, achieving an average PSNR improvement of 1.68 dB while being three times more efficient. Code will be released upon publication.
@@ 1 Introduction
Real-world images are frequently degraded by factors like haze, rain, snow, low-light, and motion blur (tian2025degradation; jiang2025survey). These spatially varying degradations may occur alternately or even overlap, depending on environmental and capture conditions (guo2024onerestore). Such entangled degradation patterns significantly reduce visual quality, resulting in impaired downstream vision tasks (tian2025degradation; jiang2025survey). Recent deep learning-based single-task methods (dong2020multi; valanarasu2022transweather; chen2021all; cai2023retinexformer; wang2022uformer; zhang2017beyond) have made significant progress in addressing individual restoration challenges. However, deploying separate task-specific (TS) networks for each degradation type is computationally expensive and impractical, driving interest in AIR frameworks (jiang2025survey).
@@
Current AIR approaches can be broadly categorized into two paradigms. (1) Adaptive feature learning methods, such as PromptIR (vaishnav2023promptir), ADAIR (cui2024adair) and AirNet (li2022all), utilize hand-crafted frequency priors or architectural inductive biases to automatically differentiate degradation types. Although these methods circumvent manual specification, they inherently depend on pre-defined assumptions about degradation characteristics, which often fail to generalize to novel or compound corruptions (gao2024prompt). (2) Prompt-based restoration methods, such as UniRestore (chen2025unirestore), InstructIR (conde2024high) and DFPIR (tian2025degradation), rely on explicit, manually provided prompts to guide the restoration network by specifying degradation types. These methods offer flexible control but face a fundamental ”chicken-and-egg” dilemma: in real-world scenarios, the degradation type and location are rarely known beforehand, yet the network requires this information to perform effective restoration (jiang2025survey). Moreover, TS instructions constrain generalization, particularly in cases of mixed degradations or varying homogeneous degradations (e.g., different noise levels). As illustrated in Fig. 1, both existing AIR paradigms have limited robustness and practical applicability in unconstrained environments.
@@
To overcome these limitations, we reframe AIR as a latent prior inference problem. Unlike existing methods (chen2025unirestore; conde2024high; tian2025degradation) that rely on explicit degradation prompting or predefined architectural inductive biases, we propose learning degradation-aware representations directly from the degraded image through a multi-level feature descriptor inspired by the variational autoencoder (VAE) (kingma2013auto). Our learned prior eliminates the need for manual degradation hints and enables spatially adaptive restoration of diverse and unseen degradations. Guided by these learned priors, the proposed degradation-aware AIR (DAIR) framework incorporates a ”where, which, what” reasoning paradigm: it learns degradation-aware feature selection (which), localizes corrupted regions using spatial attention maps (where), and adaptively fuses multi-scale global representations (what). Our unified strategy substantially enhances the flexibility and generalizability of blind restoration methods.
@@
  • DAIR consistently outperforms SOTA on six common restoration tasks (e.g., snow, low-light), five compound degradations (e.g., haze + rain, low-light + haze + rain), and unseen degradations, achieving an average 1.68 dB PSNR gain. It also improves downstream tasks on images with unknown multi-type degradations, e.g., boosting YOLOv12-L (tian2025yolov12) object detection (OD) by up to 7.40 mAP over SOTA AIR methods, highlighting strong generalizability (Fig. 1).
@@ 2.1 Image-to-image Translation
Early image restoration methods were designed to address individual TS settings, such as denoising (zhang2017beyond; pang2021recorrupted; zhang2018ffdnet), dehazing (dong2020multi; qin2020ffa), deraining(jiang2020multi; ren2019progressive), low-light image enhancement (LLIE) (wei2018deep; yi2023diff), or deblurring (cho2021rethinking; nah2017deep). These task-dependent approaches achieved impressive results; their limited scope prevented generalization across diverse restoration challenges. Transformer-based models, such as Restormer (zamir2022restormer) and Uformer (wang2022uformer), have explored multiple degradation scenarios; however, they still require separate training for individual tasks. To address this, recent methods shift towards AIR, aiming to handle diverse degradations within a unified framework. AirNet (li2022all) pioneered this field by introducing contrastive learning to extract degradation representations, guiding restoration for unknown corruptions. ADAIR (cui2024adair) recalibrate features using adaptive frequency statistics, enabling task-agnostic restoration. Recent methods, such as DiffUIR (zheng2024selective), have also explored generative techniques (i.e., latent diffusion (rombach2022ldm)). These unified approaches mark a significant step forward, minimizing TS training and advancing toward generalizable AIR solutions.
@@ 2.2 Prompt-Guided All-in-one Restoration
Image-to-image translation-based method commonly fails in separating unique degradations (brooks2023instructpix2pix; jiang2025survey). To counter this, recent AIR (gao2024prompt; conde2024high; chen2025unirestore; tian2025degradation) methods have incorporated MP to provide prior knowledge of degradation cues. InstructIR (conde2024high) and UniRestore (chen2025unirestore) integrates text-based prompts into a transformer framework, leveraging degradation-specific semantic information to guide the restoration process. OneRestore (guo2024onerestore) introduced visual and text prompts within a transformer-based framework to encode TS information. Recently, DFPIR (tian2025degradation) introduced a degradation-aware feature perturbation method that utilizes CLIP-encoded (radford2021learning) text prompts to guide channel shuffling and attention masking, enabling unified restoration across diverse degradations. UHD-Processor (liu2025uhd) introduced a VAE-based framework with progressive frequency learning and MP for ultra-high-definition image restoration. Notably, unlike UHD-Processor and other prompt-based methods, our method learns latent degradation representations directly from the input, enabling blind restoration without user guidance.
Figure 2: Overview of the proposed DAIR framework. (a) The overall architecture consists of an encoder-decoder structure complemented by degradation priors from a VAE encoder. The degraded image is encoded by our Latent Prior Encoding (LPE) in both luminance and chrominance spaces. (b) Latent Prior Encoding (which) comprises multiple Degradation Aware Encoder Blocks (DAEB) that encode degradation prior information from the VAE. Latents encoded at different stages are passed to the Degradation Map blocks. (c) The Degradation Mapping (where) block utilizes LPE latents and VAE priors to generate degradation maps for each stage. (d) The Latent Fusion (what) block combines luminance and chrominance latents to generate unified latents.
@@ 3 Method
Fig. 2 illustrates the details of our framework. Given a degraded input image $\mathbf{I}_{deg}\in\mathbb{R}^{3\times H\times W}$ with unknown degradation $\mathcal{D}_{R}$, our framework learn to infer stage-wise latent codes $x\ell$ ($\ell=0,\dots,3$) and a bottleneck global descriptor $\mu$, serving as degradation priors. The encoder employs a two-branch architecture that separately processes multi-scale structural features (e.g., luminance $L_{\ell}$) and color features (e.g., chrominance $C_{\ell}$), enabling more effective representation of both spatial and chromatic information (sharif2025illuminatingdarknesslearningenhance; yan2025hvi). At each stage, $L_{\ell}$ and $C_{\ell}$ interact with $x_{\ell}$ via Degradation-Aware Encoder Blocks (DAEB), leveraging degradation information into the feature extraction process. The branches are fused as $F_{\ell}=L_{\ell}+C_{\ell}$ and used as residual refined feature propagation to the decoder. Concurrently, a mapping block transforms $(L_{\ell},C_{\ell},x_{\ell})$ into a compact DM $g_{\ell}$, which is also forwarded to the decoder for explicit degradation localization guidance. At the bottleneck, a $\mu$-guided fusion block merges the deepest LC features with the global degradation latent using adaptive scaling and shifting (perez2018film), resulting in $F_{\mu}=\gamma\odot[L_{3}+C_{3}]+\beta$. The decoder reconstructs the clean image from $F_{\mu}$, leveraging the degradation focus map $g_{\ell}$ and the selected features $F_{\ell}$ derived from latent priors. This architecture enables the model to automatically identify and restore a wide range of degradations without explicit user input or task specification.
@@ 3.1 Learning Degradation Reasoning with Latent Prior
We propose a multi-scale latent descriptor (kingma2013auto) that directly infers continuous degradation representations from corrupted images without user-guided MP. Given $\mathbf{I}_{deg}\in\mathbb{R}^{3\times H\times W}$, the encoder extracts hierarchical feature maps $\{\mathbf{x}_{\ell}\}_{\ell=0}^{3}$ using convolutional layers and residual channel-attention (hu2018squeeze) blocks that emphasize degradation-relevant cues. At the deepest stage, multi-head self-attention (vaswani2017attention) refines the features to obtain the bottleneck representation $\mathbf{z}$. The VAE parameterizes a distribution with $\bm{\mu}=f_{\mu}(\mathbf{z})$ and $\log\bm{\sigma}^{2}=f_{\sigma}(\mathbf{z})$, sampling $\mathbf{z}_{reparam}=\bm{\mu}+\bm{\sigma}\odot\bm{\epsilon}$ where $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. The decoder reconstructs the input image from the sampled latent. This multi-objective training ensures that hierarchical features $\{\mathbf{x}_{\ell}\}$ capture degradation-specific representations at multiple scales, while bottleneck statistics $\bm{\mu}$ encode global degradation semantics. Both serve as degradation-aware latent cues to condition the restoration model:
@@ 3.1.1 Latent Prior Encoding
Unlike existing methods (DFPIR (tian2025degradation), CAPTNet (gao2024prompt)) that inject prompts during decoding, we inject VAE-derived degradation priors during encoding. This enables our encoders to generate both degradation-aware and degradation-agnostic skip connections, providing richer guidance for reconstruction. We separately encode the luminance and chrominance (LC) components (smith1978color)($\mathbf{L}_{\text{deg}}$, $\mathbf{C}_{\text{deg}}$) of the degraded input using dedicated LC encoders (yan2025hvi; sharif2025illuminatingdarknesslearningenhance) for better structural and chromatic information. Fig. 2 (b) illustrates our latent prior encoding. Two parallel encoders, $\mathcal{E}_{L}(\cdot)$ and $\mathcal{E}_{C}(\cdot)$, process these components across multiple scales using DAEB. At stage $\ell$, we injects latent degradation priors $\mathbf{x}_{\ell}$ via value modulation:
@@ 3.1.2 Learnable Degradation Map
Traditional skip connections propagate noise from degraded inputs to the decoder, without spatial awareness of the degradation (mao2016image). We counter these limitations with learnable DM that encode spatially-varying degradation patterns for effective reconstruction (Fig. 2 (c)). At encoder stage $\ell$, our Degradation Mapping Block generates scale-specific maps $g_{\ell}$ from LC features $L_{\ell},C_{\ell}$ and latent prior encoding $\mathbf{x}_{\ell}$. We first fuse complementary LC information: $\mathbf{f}^{LC}_{\ell}=\text{Conv}_{3\times 3}(L_{\ell}+C_{\ell})$. To capture frequency-domain degradation characteristics, we enhance features with spectral information via Fast Fourier Transform (cooley1965algorithm): $\mathcal{F}(L_{\ell})=\text{FFT}(L_{\ell}-\mu(L_{\ell}))$, $\mathcal{F}(C_{\ell})=\text{FFT}(C_{\ell}-\mu(C_{\ell}))$. The magnitude and phase components are concatenated and projected:
@@ 3.1.3 Latent Encoded Fusion
At the encoder bottleneck, we leverage the deepest LC features $\mathbf{L}_{3},\mathbf{C}_{3}\in\mathbb{R}^{B\times 320\times\frac{H}{8}\times\frac{W}{8}}$, along with the latent global descriptor $\mu$. As shown in Fig. 2 (d), our fusion mechanism integrates three components: (i) intra-branch self-attention (sharif2025illuminatingdarknesslearningenhance; cai2023retinexformer) for independent modality refinement, where $\hat{\mathbf{L}}_{3}=\text{MHSA}(\mathbf{L}_{3})$ and $\hat{\mathbf{C}}_{3}=\text{MHSA}(\mathbf{C}_{3})$; (ii) cross-branch attention (sharif2025illuminatingdarknesslearningenhance) for complementary information exchange, where chrominance attends to luminance: $\mathbf{F}=\text{LayerNorm}\left(\text{MHCA}(\hat{\mathbf{C}}_{3},\hat{\mathbf{L}}_{3})+\hat{\mathbf{C}}_{3}+\hat{\mathbf{L}}_{3}\right)$, enabling the model to leverage structural and color information while preserving individual branch characteristics; and (iii) adaptive feature modulation conditioned on the global degradation prior $\mu$, combining LC features through adaptive scaling and shifting for TS feature modulation while maintaining underlying image characteristics. Unlike standard FiLM (perez2018film) that applies uniform modulation, our approach makes degradation-aware reconstruction: